Working capital: Laurel Electronics reported the following information at its annual meetings: The company had cash and marketable securities worth 1,305,449, accounts payables worth 4,188,061, inventory of 7,893,603, accounts receivables of 3,458,080, short-term notes payable worth 1,224,947, and other current assets of 123,457. What is the company’s net working capital? Round your answer to a whole number.
Answer #1

Total current assets=cash and marketable securities+accounts receivable+inventory+other current assets

=1,305,449+7,893,603+3,458,080+123,457=$12780589

Total current liabilities=accounts payable+short term notes payable

=4,188,061+1,224,947=$5413008

Net working capital=Total current assets-Total current liabilities

=12780589-5413008

=$7367581